OpenStudy (anonymous):

An object is thrown upward from the edge of a tall building with a velocity of 10 m/s. Where will the object be 3 s after it is thrown? Take

OpenStudy (hunus):

Do you have the height of the building?

OpenStudy (shamim):

the formula is\[v=u- g t\]

OpenStudy (shamim):

for maximum height final velocity v=0

OpenStudy (shamim):

\[0=10- 10 t\]

OpenStudy (shamim):

t=1 sec

OpenStudy (shamim):

so time required to reach its maximum height is 1 sec

OpenStudy (shamim):

it will come back to the top of the building after another 1 sec

OpenStudy (shamim):

so total

OpenStudy (shamim):

time needed to come back to the top of the building is 2 sec

OpenStudy (shamim):

ur throughing object will fall more in another 1 sec if the building is very high

OpenStudy (anonymous):

it will be 20m (9.8*2), after reaching its max height (t=1)

OpenStudy (anonymous):

20m/s

OpenStudy (anonymous):

correct shamim?

OpenStudy (shamim):

no

OpenStudy (shamim):

u know the equation for maximum height is\[v ^{2}=u ^{2}- 2 g h\]

OpenStudy (shamim):

u know for maximum height final velocity v=0

OpenStudy (shamim):

so the equation will b\[0^{2}=10^{2}-2 \times 10 \times h\]

OpenStudy (shamim):

anyway i m taking gravitational acceleration g=10m/s

OpenStudy (shamim):

h=5m

OpenStudy (shamim):

maximum height h=5m

OpenStudy (shamim):

ya u r right @dgamma3

OpenStudy (shamim):

u were telling the final velocity after 3 sec

OpenStudy (shamim):

understood

OpenStudy (mos1635):

there is an other way (equation) s=Vo*t -1/2*g*t^2 Vo=10 t=3 g=10 (9.8) solve s= -15m (that means 15 m below point of throw)

OpenStudy (shamim):

ya u r right @mos